OpenStudy (flvskidd):

Max and Maggie have to clean the house. It takes Max 12 hours to clean the house, while Maggie can complete the task in 4 hours. Their sister says that it will take 3 hours to complete if they work together. Explain each step in solving this equation, and determine if the sister is correct or not.

OpenStudy (anonymous):

you want the quick answer?

OpenStudy (flvskidd):

i think my teacher would want an explanation

OpenStudy (anonymous):

Good answer kid.

OpenStudy (anonymous):

using only the numbers given, 12 and 4, it will take them \[\frac{12\times 4}{12+4}\] hours to do it if you really want to impress your teacher, show where this easy answer comes from

OpenStudy (anonymous):

it is the reciprocal of what you get when you add the fraction one has rate \(\frac{1}{12}\) the other has rate \(\frac{1}{4}\) their combined rates is what you get when you add them, namely \[\frac{1}{12}+\frac{1}{4}=\frac{12+4}{12\times 4}\]

OpenStudy (anonymous):

you then want to see how long it takes to complete ONE job (this is usually the confusing part) , in other words solve \[\frac{12+4}{12\times 4}T=1\] for \(T\) btw you can compute that number at any stime

OpenStudy (anonymous):

solve for \(T\) via \[T=\frac{12\times 4}{12+4}\]

OpenStudy (flvskidd):

T = 48/16

OpenStudy (anonymous):

reduce to make it look nicer